# The Frequency Response of Networks as Open Systems

### v1

@@ II.5 Directed Acyclic Graphs
We focus on open networks defined by the triplet $(A,B,C)$, where $A=P-D$. Here, $P=[P_{ij}]$, with $P_{ij}\geq 0$, represents the adjacency matrix of a DAG, and $D$ is a diagonal matrix with non-negative entries. By construction, the matrix $A$ is a Metzler matrix. The quantity of interest is the $ji$-th entry of $A^{-1}$, which corresponds to the DC gain when node $i$ is chosen as the input and node $j$ as the output. Following [63, Theorem 6.4], it is easy to calculate
$[A^{-1}]_{ji}=\sum_{\text{directed path }\pi:i\to j}\left(\prod_{k\to l\in\pi}\dfrac{A_{lk}}{-A_{kk}}\right)\dfrac{1}{A_{jj}}.$ (13)
@@
$[A^{-1}]_{ji}=-\sum_{\text{directed path }\pi:i\to j}\left(\prod_{k\to l\in\pi}\dfrac{A_{lk}}{|A_{kk}|}\right)\dfrac{1}{|A_{jj}|}.$ (14)
It is important to highlight the intuitive interpretation of Eq. (14). We see that the passing, amplifying, or damping behavior of DAG networks as open systems depends both on the number of directed paths from input to output and on the weights of the edges that make up these paths. Each edge weight contributes to amplification if it is larger than one, or to attenuation if it is smaller than one; longer paths compound these effects. Moreover, the self-loops always contribute to damping—since in our formulation they correspond to the local stability term $c$—and appear in the denominator of each term in the sum. Based on this reasoning, we can also understand why gene networks, which have many but very short (often one-step) paths, exhibit a considerable decrease in the normalized trace $\text{Tr}(W_{c})/N^{2}$, whereas food webs and signaling pathways, with fewer but longer weighted paths, maintain higher normalized trace values.
@@
DC $\text{DC Gain}=20\log_{10}\left|[A^{-1}]_{ji}\right|,$ (15)
where $[A^{-1}]_{ji}$ is given explicitly in Eq. (14). Next we present an illustrative example.
@@
Figure  5 shows a directed acyclic graph (without considering self-loops) and its adjacency matrix $A$. For this network, node $i=1$ is the input node and node $j=5$ is the output node. There are $3$ paths from node 1 to 5: $\pi_{1}=(1,2,4,5)$, $\pi_{2}=(1,2,3,4,5)$, and $\pi_{3}=(1,3,4,5)$. Therefore, applying Eq. (14) results in
@@
FIG. 5: Directed Acyclic Graph (excluding self-loops) and its corresponding adjacency matrix $A$. The input and the output nodes are shown in blue and orange, respectively. Self-loops are considered to provide asymptotic stability of the linear dynamics.
@@ II.6 Frequency Response of General Networks
Next, we consider general network topologies (not necessarily having a DAG structure), under the only assumption that the weights of the connections (excluding self-loops) are non-negative. This assumption is satisfied by all data sets in this study. Our numerical results indicate that the slope of the magnitude plot for large frequencies ($\omega\gg$ cornering frequency) and the final phase for large frequencies follows the general relations,
@@
where $d$ is the shortest path length from the input node to the output node without considering the weights. Alternatively, $d+1$ is the effective order of the dominant pole at the cornering frequency $c$, where $-c$ is the largest real-part eigenvalue of the matrix $A$. Figure 6 confirms our above relations for the slope, the final phase, and the cornering frequency. Note that in this example, if one considers the weights when evaluating $d$, then the path from the input node to node 2 to the output node is no longer the shortest, while the shortest path involves visiting node 4 before node 2. The evaluation of the weighted shortest path does not match the numerical results shown in the left panels of Fig. 6, while our definition of the shortest unweighted path length matches.
FIG. 6: Bode diagram for a general weighted network. Bode plot (left panels) of the network shown on the right. The matrix $A$ defining the topology of the network is shifted as $A\leftarrow A-cI$ such that the largest real part eigenvalue of the matrix $A$ is $-1$. All nodes have self-loops of weight $-c$, which are not shown for better visualization. The shortest path without considering edge weights is from the input node to node 2 and then to the output node, which has a length of $d=2$. The slope of the magnitude after the cornering frequency $c=1$ is slope $=-60$, and the final phase $=-270$.
@@ III Conclusion
In this work, we propose a general theory to study how networked dynamical systems behave as open systems—processing, amplifying, or attenuating environmental inputs depending on their internal structure and the specific configuration of input and output nodes. We employ the $\mathcal{H}_{2}$-norm as a robust and universal metric to quantify the degree of input-to-output signal amplification across a wide class of inputs, including sinusoidal, periodic, and stochastic signals.
@@
Our theoretical results demonstrate that the signal amplification characteristics of a network are strongly influenced by structural factors such as the location of input and output nodes, the node-to-node weighted connectivity, and unweighted shortest path distances. We established relationships between the network topology and signal transmission behavior, providing conditions under which a network tends to either pass or block signals, based on frequency response and DC gain analyses. Importantly, we identified network topologies that exhibit optimized configurations, either facilitating transmission (passing) or suppressing disturbances (blocking), reflecting functional demands of the network with its environment and constraints arising from evolution or design.
@@
An empirical analysis of networks across diverse domains, including power grids, gene regulatory systems, and neural connectomes, support our theoretical predictions and highlight the broad applicability of our framework. We find that food webs, signaling pathways, and gene regulatory circuits are often structurally organized to enhance the passing of signals, while the structure of engineered systems like power grids appears to be intentionally designed to suppress signal and noise propagation. Overall, this study lays the groundwork for systematic characterization and optimization of open networks, with potential applications in network design, control, and the interpretation of naturally occurring network dynamics.
@@
Most of previous effort in investigating the response of network-coupled dynamical systems to external inputs has focused on undirected networks. Many of the results obtained for undirected networks do not generalize to directed ones, which are typically more challenging to investigate. In this work, we have applied tools and obtained results that apply to generic (non-normal) network topologies and uncovered that directedness can be used as a fundamental passing mechanism. In particular, our data analysis reveals that directed acyclic graphs (DAGs)—a topological feature prevalent in food webs, signaling pathways, and gene regulatory networks—exhibit optimal passing behavior. This empirical observation aligns with our analytical results for DAGs, which demonstrate that amplification and attenuation depend explicitly on the number and length of the weighted input–output paths, as well as on attenuation effects at the node level.
@@
Our work provides direct insight into how to tune the $\mathcal{H}_{2}$-norm of an open network—either increasing it by strategically adding input or output nodes, or reducing it by removing them. This insight enables a principled approach to ranking nodes based on their contribution to the network’s amplification properties, leveraging the modularity of the output controllability Gramian. These findings have direct implications for engineering and bioengineering applications, for which achieving an optimal balance between blocking and passing behavior is often critical.
@@ IV Methods
IV.1 DC Gain
@@
In this limit of the frequency variable approaching zero, $\omega\to 0$, the transfer function $G(J\omega)$ simplifies as follows:
DC $\begin{split}\text{DC Gain}&=\lim_{\omega\to 0}20\log_{10}|G(J\omega)|\\ &=\lim_{\omega\to 0}20\log_{10}|C(J\omega-A)^{-1}B|\\ &=20\log_{10}|CA^{-1}B|,\end{split}$
$\begin{split}\text{Phase}=\lim_{\omega\to 0}\angle G(J\omega)=\lim_{\omega\to 0}\angle C(J\omega-A)^{-1}B\\ =\angle(-CA^{-1}B).\end{split}$
@@
In the case where a control signal is applied to node $i$ and the response is measured at node $j$, so that $B=\boldsymbol{e}_{i}$, $C=\boldsymbol{e}_{j}^{\top}$, the expressions simplify to
DC $\text{DC Gain}=20\log_{10}\left|[A^{-1}]_{ji}\right|=20\log_{10}\left|\frac{Q_{ij}}{\det(A)}\right|,$ (17)
@@
IV.2 Properties of Controllability Gramian
Controllability Gramian is a symmetric and positive semi-definite matrix. The system is called controllable if and only if the matrix is nonsingular. Controllability Gramian can be directly related to the energy required to steer the system around the state space. Assuming the initial condition for the system is at the origin and the desired final state is $\boldsymbol{x}_{f}$, then the solution to the minimum energy problem is

### v2

@@ II.5 Directed Acyclic Graphs
We focus on open networks defined by the triplet $(A,B,C)$. The off-diagonal entries of the matrix $A$ represent the connectivity of a DAG, while the diagonal entries are negative quantities that represent the stabilizing self feedback coefficient at each individual node. By construction, here, we take the matrix $A$ to be a Metzler matrix, i.e., a matrix for which all the off-diagonal entries are nonnegative. The quantity of interest is the $ji$-th entry of $A^{-1}$, which corresponds to the DC gain when $i$ is chosen as the input node and $j$ as the output node. Following [69, Theorem 6.4], it is easy to calculate
$A^{-1}]_{ji}=\sum_{\text{directed path }\pi:i\to j}\left(\prod_{k\to l\in\pi}\dfrac{A_{lk}}{-A_{kk}}\right)\dfrac{1}{A_{jj}}.$ (13)
@@
$A^{-1}]_{ji}=-\sum_{\text{directed path }\pi:i\to j}\left(\prod_{k\to l\in\pi}\dfrac{A_{lk}}{|A_{kk}|}\right)\dfrac{1}{|A_{jj}|}.$ (14)
It is important to highlight the intuitive interpretation of Eq. (14). We see that the passing, amplifying, or damping behavior of DAG networks as open systems depends both on the number of directed paths from input to output and on the weights of the edges that make up these paths. Each edge weight contributes to amplification if it is larger than one, or to attenuation if it is smaller than one; longer paths compound these effects. Moreover, the self-loops always contribute to damping—since in our formulation they correspond to the local stability term cc—and appear in the denominator of each term in the sum. Based on this reasoning, we can also understand why gene networks, which have many but very short (often one-step) paths, exhibit a considerable decrease in the normalized trace $\text{Tr}(W_{c})/N^{2}$, whereas food webs and signaling pathways, with fewer but longer weighted paths, maintain higher normalized trace values.
@@
$\text{DC Gain}=20\log_{10}\left|[A^{-1}]_{ji}\right|,$ (15)
where $A^{-1}]_{ji}$ is given explicitly in Eq. (14). Next we present an illustrative example.
@@
Figure  5 shows a directed acyclic graph (without considering self-loops) and its associated matrix $A$. For this network, node $i=1$ is the input node and node $j=5$ is the output node. There are $3$ paths from node 1 to 5: $\pi_{1}=(1,2,4,5)$, $\pi_{2}=(1,2,3,4,5)$, and $\pi_{3}=(1,3,4,5)$. Therefore, applying Eq. (14) results in
@@
FIG. 5: Directed Acyclic Graph (excluding self-loops) and its corresponding Jacobian matrix $A$. The input and the output nodes are shown in black and orange, respectively. Self-loops are considered to provide asymptotic stability of the linear dynamics.
@@ II.6 Frequency Response of General Networks
Next, we consider general network topologies—without requiring a DAG structure—while retaining the assumption that the matrix $A$ is Metzler and has negative real part eigenvalues. We find that the slope of the magnitude plot for large frequencies ($\omega\gg$ cornering frequency) and the final phase for large frequencies follows the general relations,
@@
where $d$ is the shortest path length from the input node to the output node without considering the weights. Alternatively, $d+1$ is the effective order of the dominant pole at the cornering frequency $c$, where c-c is the largest real-part eigenvalue of the matrix $A$. Figure 6 confirms our above relations for the slope, the final phase, and the cornering frequency. Note that in this example, if one considers the weights when evaluating $d$, then the path from the input node to node 2 to the output node is no longer the shortest, while the shortest path involves visiting node 4 before node 2. The evaluation of the weighted shortest path does not match the numerical results shown in the left panels of Fig. 6, while our definition of the shortest unweighted path length matches.
FIG. 6: Bode diagram for a general weighted network. Bode plot (left panels) of the network shown on the right. The matrix AA defining the topology of the network is shifted as $A\leftarrow A-cI$ such that the largest real part eigenvalue of the matrix $A$ is 1-1. All nodes have self-loops of weight c-c, which are not shown for better visualization. The shortest path without considering edge weights is from the input node to node 2 and then to the output node, which has a length of $d=2$. The slope of the magnitude after the cornering frequency $c=1$ is slope =60=-60, and the final phase =270=-270.
@@ III Conclusion
In this work, we propose a general theory to study how networked dynamical systems behave as open systems—processing, amplifying, or attenuating environmental inputs depending on their internal structure and the specific configuration of input and output nodes. We employ the $\mathcal{H}_{2}$-norm as a robust and universal metric to quantify the degree of input-to-output signal amplification across a wide class of inputs, including sinusoidal, periodic, and stochastic signals. Within this framework, we show how the frequency response of an arbitrary network, including its Bode profile, is affected jointly by the underlying topology and by the choice of input and output nodes. We also examine the $\mathcal{H}_{2}$-norm across several empirical networks and derive new closed-form results for directed acyclic graphs and for the zero-frequency limit. In addition, we obtain an asymptotic approximation for the trace of the Gramian in the regime dominated by strongly stable dynamics, and we uncover a modularity property of the squared $\mathcal{H}_{2}$-norm that applies to both input and output nodes, extending previous results that focused only on inputs.
@@
Our theoretical results demonstrate that the signal amplification characteristics of a network are strongly influenced by structural factors such as the location of input and output nodes, the node-to-node weighted connectivity, and unweighted shortest path distances. We establish relationships between the network topology and signal transmission behavior, providing conditions under which a network tends to either pass or block signals, based on frequency response and DC gain analyses. Importantly, we identify network topologies that exhibit optimized configurations, either facilitating transmission (passing) or suppressing disturbances (blocking), reflecting functional demands of the network with its environment and constraints arising from evolution or design. The work in [70] investigates open networks in discrete time; however, it does not consider their frequency response.
@@
An empirical analysis of networks across diverse domains, including power grids, gene regulatory systems, and neural connectomes, supports our theoretical predictions and highlights the broad applicability of our framework. We find that food webs, signaling pathways, and gene regulatory circuits are often structurally organized to enhance the passing of signals, while the structure of engineered systems like power grids tends to suppress signal and noise propagation. Overall, this study lays the groundwork for systematic characterization and optimization of open networks, with potential applications in network design, control, and the interpretation of naturally occurring network dynamics. This exploration of empirical network structures, together with the dynamics they are supporting complement and further previous works which focused on the effect of topology and nonlinearities on the propagation of perturbations or on specific network structures that amplify or reduce fluctuations [26, 27, 71].
@@
Most of previous effort in investigating the response of network-coupled dynamical systems to external inputs has focused on undirected networks. Many of the results obtained for undirected networks do not generalize to directed ones, which are typically more challenging to investigate. In this work, we have applied tools and obtained results that apply to generic (non-normal) network topologies and uncovered that directedness can be used as a fundamental passing mechanism. In particular, our data analysis reveals that directed acyclic graphs (DAGs)—a topological feature that is prevalent in food webs, signaling pathways, and gene regulatory networks—exhibit strong passing behavior. This empirical observation aligns with our analytical results for DAGs, which demonstrate that amplification and attenuation depend explicitly on the number and length of the weighted input–output paths, as well as on attenuation effects at the node level. Taken together, these findings are consistent with the previous studies of directed, non-normal networks, where strong asymmetries in the interaction pattern generate source–sink hierarchies, suppress back-propagation, and bias flows directionality. In such architectures, which are often structurally close to acyclic and possess a dominant directed backbone with only weak feedback loops, perturbations are funneled toward terminal components and sink nodes, naturally favoring one-way signal transfer [17, 18, 19, 22, 25].
While our analysis captures how network structure shapes signal propagation, it relies on the linear dynamics encoded in the Jacobian of the underlying nonlinear system. As we now emphasize, the LTI description reflects only the first–order behavior around a fixed point, and different nonlinear models defined on the same structure—or even different equilibria of the same model—can yield different Jacobians and thus different response patterns. This is an inherent limitation of any linearization-based approach. Nevertheless, linear analysis often provides meaningful insight into regimes where systems operate near criticality or in weakly nonlinear phases, where the Jacobian accurately captures local signal propagation. Outside these conditions, fully nonlinear effects become essential and lie beyond the scope of the present work. We also note that some of our results, in particular those in Secs. II E and II F, are limited to the case of non-negative weights associated with the network edges.
@@
Our work provides direct insight into how to tune the $\mathcal{H}_{2}$-norm of an open network—either increasing it by strategically adding input or output nodes, or reducing it by removing them. This insight enables a principled approach to ranking nodes based on their contribution to the network’s amplification properties, leveraging the modularity of the output Gramian. These findings have direct implications for engineering and bioengineering applications, for which achieving an optimal balance between blocking and passing behavior is often essential.
Several key questions about how complex systems respond to perturbations remain open. For instance, time-varying coupling—arising from failures or from the intrinsic evolution of the underlying network—may strongly affect signal transmission, and the impact of control actions triggered by propagating perturbations is still poorly understood. Furthermore, mounting evidence points to the pivotal role of higher-order (beyond pairwise) interactions in networked systems [72, 73, 74, 75, 76, 77], raising the question of how such multi-body couplings modulate their response and resilience to perturbations. Addressing these issues represents an important direction for future work.
@@ IV Methods
IV.1 From Nonlinear to Linearized equations
We start from the general set of equations for the network dynamics we consider in this paper,
$\dot{z}_{i}(t)=f(z_{i}(t))+\sigma\sum_{j}\tilde{A}_{ij}\,g(z_{i}(t),z_{j}(t)),$ (17)
$i=1,..,N$, where $z_{i}(t)\in\mathbb{R}$ is the state of system $i$, the function $f:\mathbb{R}\rightarrow\mathbb{R}$ describes the intrinsic dynamics of each node, $g:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ the pairwise interactions, the scalar $\sigma$ the global coupling strength, and $\tilde{A}\in\mathbb{R}^{N\times N}$ the directed adjacency matrix of the network. The specific forms of the functions $f$ and $g$ depend on the underlying network dynamics—for example, they differ for food webs, power grids, and connectomes—as detailed in the Methods Section IV.5.
Linearization of Eqs. (17) around a stable fixed point $z_{i}^{*}$, i=1,,Ni=1,...,N, yields the LTI dynamics,
$\dot{x}_{i}(t)=A_{ii}\,x_{i}(t)+\sum_{j\neq i}A_{ij}\,x_{j}(t),$ (18)
i=1,,Ni=1,...,N, with $A_{ii}=f^{\prime}(z_{i}^{*})+\sigma\sum_{j}\tilde{A}_{ij}\,\partial_{1}g(z_{i}^{*},z_{j}^{*})$ capturing the local stability of the intrinsic dynamics together with the self-dependence of the interactions, and $A_{ij}=\sigma\tilde{A}_{ij}\,\partial_{2}g(z_{i}^{*},z_{j}^{*})$ encoding the linearized influence of node $j$ on node $i$. Note that this equation is equivalent to our main equation (1) in the absence of inputs $u(t)=0$.
As an example, we consider an $N=6$-node network with adjacency matrix,
$\tilde{A}=\begin{bmatrix}0&1&0&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&0&1&0\\ 0&0&1&0&1&0\\ 0&0&0&0&0&1\\ 0&0&1&0&0&0\end{bmatrix}$.
With this choice of the adjacency matrix, we adopt the Michaelis-Menten model [78], a standard framework for describing the dynamics of gene-regulatory and metabolic networks,
$\dfrac{dz_{i}(t)}{dt}=-\alpha_{i}z_{i}^{a}(t)+\kappa\sum_{j=1}^{N}\tilde{A}_{ij}\dfrac{z_{j}^{h}(t)}{1+z_{j}^{h}(t)},$ (19)
$i=1,\ldots,N$, where $z_{i}(t)$ represents the expression of gene $i$ at time $t$ and wthout loss of generality, we set $a=2$, $h=1$, $\kappa=1$, and $\alpha_{i}=1$ at all nodes. Further details on this model can be found in the Methods IV.5.
We compute the Jacobian about the fixed point $\boldsymbol{z}^{*}=[0.692,0.921,0.618,0.874,0.618,0.618]^{T}$,
${A}=\begin{bmatrix}-1.384&0.271&0&0&0&0\\ 0&-1.842&0.382&0.2847&0&0\\ 0&0&-1.236&0&0.382&0\\ 0&0&0.382&-1.748&0.382&0\\ 0&0&0&0&-1.236&0.382\\ 0&0&0.382&0&0&-1.236\end{bmatrix}$.
We note the similarity in the structure of the matrices $A$ and $\tilde{A}$ above. In particular, the Jacobian $A$ corresponds to a weighted version of the network adjacency matrix $\tilde{A}$ plus a diagonal matrix, since the off-diagonal entries of the Jacobian multiply the original entries of the adjacency matrix, and self loops might exist. When the fixed point is homogeneous across nodes, these weights reduce to a uniform rescaling of the adjacency matrix. Application of a spectral shift $A\leftarrow A-cI$, $c>0$, only adds a uniform damping term and does not affect the eigenvectors of the matrix $A$. As a result, the network structure and qualitative dynamical response are preserved, with the shift modifying only the global timescale [7, 8, 9].
IV.2 Frequency Response Example
Figure 7 shows an input signal that is composed of two sinusoidal signals with two different frequencies and two different amplitudes. The input signal enters the network through the black node and results in the output signal through the red node. The output signal has the same two frequencies as the input signal, but with different amplitudes and phases. We see that the amplitude of the low-frequency input is slightly amplified (from $1\to 1.4$), while the amplitude of the high-frequency input is damped (from $0.2\to 0.002$).
Technical details:
The pair $(A,B)$ is selected from our dataset for the connectome network “Macaque 30”. The network has 30 nodes and has 7 input nodes. Node 4 was selected as the input (shown as black in Fig. 7) and node 1 was selected as the output (shown as red in Fig. 7). The distance from the input to the output node is $d=2$. The input and the output signals are composed of the following components:
$\text{Low Frequency input}=\sin(2\pi t),$ (20a)
$\text{High Frequency input}=0.2\sin(40\pi t),$ (20b)
$\text{Low Frequency output}=A_{1}\sin(2\pi t+\phi_{1}),$ (20c)
$\text{High Frequency output}=A_{2}\sin(40\pi t+\phi_{2}),$ (20d)
where
$A_{1}=|G(2\pi\jmath)|=1.41,\quad\phi_{1}=\angle G(2\pi\jmath)=-86\degree,$ (21a)
$A_{2}=|G(40\pi\jmath)|=0.002,\quad\phi_{2}=\angle G(40\pi\jmath)=-251\degree,$ (21b)
where $\jmath=\sqrt{-1}$. The amplitude and the phases of the output components can be seen in the Bode plot in the bottom panel of Fig. 7 (shown as green circles for the low-frequency signal and purple squares for the high-frequency signal). Note that magnitudes of $1.41$ and $0.002$ correspond to dB values $20\log_{10}(1.41)=3$ and $20\log_{10}(0.002)=-54$, respectively.
FIG. 7: The top panel shows the passing behavior of a low-frequency input signal and blocking a high-frequency one. The bottom panel shows the Bode plot associated with the system.
IV.3 DC Gain
@@
In this limit of the frequency variable approaching zero, $\omega\to 0$, the transfer function $G(\jmath\omega)$ simplifies as follows:
$\begin{split}\text{DC Gain}&=\lim_{\omega\to 0}20\log_{10}|G(\jmath\omega)|\\ &=\lim_{\omega\to 0}20\log_{10}|C(\jmath\omega-A)^{-1}B|\\ &=20\log_{10}|CA^{-1}B|,\end{split}$
$\begin{split}\text{Phase}=\lim_{\omega\to 0}\angle G(\jmath\omega)=\lim_{\omega\to 0}\angle C(\jmath\omega-A)^{-1}B\\ =\angle(-CA^{-1}B).\end{split}$
@@
In the case where an external signal is applied to node $i$ and the response is measured at node $j$, so that $B=\boldsymbol{e}_{i}$, $C=\boldsymbol{e}_{j}^{\top}$, the expressions simplify to
$\text{DC Gain}=20\log_{10}\left|[A^{-1}]_{ji}\right|=20\log_{10}\left|\frac{Q_{ij}}{\det(A)}\right|,$ (22)
@@
IV.4 Properties of Controllability Gramian
The controllability Gramian is a symmetric and positive semi-definite matrix. The system is called controllable if and only if the matrix is nonsingular. The Gramian can be directly related to the energy required to steer the system around the state space. Assuming the initial condition for the system is at the origin and the desired final state is $\boldsymbol{x}_{f}$, then the solution to the minimum energy problem is